### v1

# Bringing closure to FDR control: beating the e-Benjamini-Hochberg procedure

Abstract
False discovery rate (FDR) has been a key metric for error control in multiple hypothesis testing, and many methods have developed for FDR control across a diverse cross-section of settings and applications. We develop a closure principle for all FDR controlling procedures, i.e., we provide a characterization based on e-values for all admissible FDR controlling procedures. We leverage this idea to formulate the closed eBH procedure, a (usually strict) improvement over the eBH procedure for FDR control when provided with e-values. We demonstrate the practical performance of closed eBH in simulations.
1 Introduction
Multiple testing with p-values, in particular false discovery rate (FDR) control, has been a centerpiece of statistical research in the past three decades since its introduction by Benjamini and Hochberg (1995). Recently, Wang and Ramdas (2022) formulated the e-Benjamini-Hochberg (eBH) procedure for FDR control, which uses e-values for each hypothesis and is robust to arbitrary dependence. It is known that by employing external randomization, or additional information about distribution of the e-values or their dependence structure, the eBH procedure can be improved (i.e. it can be amended to be more powerful).
In this paper, we will show that one can (often strictly) improve the eBH procedure with no additional information or randomization. We call our procedure the closed eBH procedure, denoted ${\overline{\text{eBH}}}$ for brevity. Further, we establish that any FDR controlling procedure must rely on e-values, and in particular can be written as an instance of ${\overline{\text{eBH}}}$ acting on some e-values. In the process, we establish a closure principle for FDR control, similar to the results of Marcus et al. (1976) and Sonnemann and Finner (1988) for FWER control. This is also complementary to the results of Ignatiadis et al. (2025) who show that every FDR controlling procedure is an instance of the eBH procedure with compound e-values, a fact that we return to later in this work.
Problem setup.
The standard multiple testing problem assumes that there are $K$ null hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{K}$ a scientist wishes to test, some of which are truly null, and one wishes to discover the non-nulls. Let $\alpha\in[0,1]$ be our fixed and desired level of control on the FDR.
First, for any set $A\subseteq[K]\coloneqq\{1,\dots,K\}$, define $\mathcal{H}_{A}\coloneqq\cap_{j\in A}\mathcal{H}_{j}$, i.e., the intersection hypotheses of the base hypotheses corresponding to the set $A$. For any set $A$, if the true null set equals $A$, then the false discovery proportion (FDP) of a candidate discovery set $R\subseteq[K]$ equals
$\text{FDP}_{A}(R)=\sum_{j\in A}\frac{\mathbf{1}\{j\in R\}}{|R|\vee 1}=\frac{|R\cap A|}{|R|\vee 1}.$
In words, if $\mathcal{H}_{A}$ happened to be true (and $\mathcal{H}_{B}$ is false for every $B\supset A$), and $R$ was rejected, then $\text{FDP}_{A}(R)$ would be the realized FDP.
Let $\mathbf{X}\in\mathcal{X}$ denote the data, which is sampled from the true distribution $\mathcal{P}^{*}$. We define a discovery procedure $\mathcal{R}:\mathcal{X}\rightarrow 2^{[K]}$ as a mapping from input data to a discovery set $R\subseteq[K]$, where $\mathbb{R}^{\geq 0}$ represents the nonnegative reals. Formally, each base hypothesis $\mathcal{H}_{j}$ is a set of distributions, and $\mathcal{H}_{j}$ is said to be “truly null” if $\mathcal{P}^{*}\in\mathcal{H}_{j}$. Let $A^{*}\subseteq[K]$ denote the set of true nulls, i.e., $\mathcal{P}^{*}\in\mathcal{H}_{A^{*}}$ and $\mathcal{P}^{*}\not\in\mathcal{H}_{A}$ for all $A$ s.t. $A^{*}\subset A$. The FDR of a discovery procedure $\mathcal{R}$ equals
$\textnormal{FDR}(\mathcal{R})\coloneqq\mathbb{E}[\textnormal{FDP}_{A^{*}}(\mathcal{R}(\mathbf{X}))].$
The eBH procedure.
We consider the situation where our data has already been summarized into $K$ e-values $\mathbf{X}=\mathbf{E}\coloneqq(E_{1},\dots,E_{K})$ (one per hypothesis). These e-values may be arbitrarily dependent. Wang and Ramdas (2022) introduced the eBH procedure which takes in ensures FDR control below $\alpha$ under arbitrary dependence among the e-values, and is defined as producing the following discovery set:
$k^{\textnormal{eBH}}$ $\coloneqq\max\ \left\{k\in[K]:\sum_{i\in[K]}\mathbf{1}\left\{E_{i}\geq\frac{K}{\alpha k}\right\}\geq k\right\}$ (1)
$R^{\textnormal{eBH}}$ $\coloneqq\mathcal{R}^{\textnormal{eBH}}(\mathbf{E})\coloneqq\left\{i\in[K]:E_{i}\geq\frac{K}{\alpha k^{\textnormal{eBH}}}\right\}$ (2)
The eBH procedure (and variants) has been used to derive FDR control in many settings with complex dependencies in the data, e.g., bandit data collection (Xu et al., 2021), multi-stream monitoring (Wang et al., 2025; Dandapanthula and Ramdas, 2025; Tavyrikov et al., 2025), multi-resolution testing (Gablenz and Sabatti, 2025; Yu et al., 2025), combining data from multiple sources (Banerjee et al., 2024; Li and Zhang, 2025), etc. Prior work has shown that existing FDR controlling methodologies can be recast as an application of the eBH procedure. Examples include knockoffs (Ren and Barber, 2023; Guan et al., 2025) as well as the seminal procedure of Benjamini and Hochberg (1995) (Li and Zhang, 2025) and its variant introduced by Benjamini and Yekutieli (2001) that is robust to arbitrary dependence (Wang and Ramdas, 2022). We will show that the eBH procedure is actually inadmissible, and derive a procedure that (usually strictly) improves upon eBH that is applicable to and may improve power for all of the aforementioned situations.
Paper outline.
We introduce the ${\overline{\text{eBH}}}$ procedure in Section 2. Then, we formulate our closue principle for all FDR controlling procedures in Section 3. In Section 4, we formulate the closed compound eBH procedure for compound e-values. We describe some extensions and synergies of the ${\overline{\text{eBH}}}$ procedure with existing e-value techniques (i.e., randomization, post-hoc validity, e-merging, boosting). We present simulations showing the improvements of ${\overline{\text{eBH}}}$ over existing eBH methodology in Section 6. Lastly, we discuss prior work in depth in Section 7 and recap our findings in Section 8.
2 The closed eBH (${\overline{\text{eBH}}}$) procedure
Our procedure rejects more hypotheses (often strictly) than eBH, while still controlling the FDR. The construction proceeds akin to the closure principle for familywise error control (Marcus et al., 1976) by first defining a test for every intersection hypothesis. Define the e-value to test $\mathcal{H}_{A}$ as
$E_{A}=|A|^{-1}\sum_{i\in A}E_{i}.$
The arithmetic mean is a symmetric admissible e-value merging function (Vovk and Wang, 2021). Let $\mathbb{E}_{A}$ denote the supremum of the expectations of a random variable over distributions where $\mathcal{H}_{A}$ holds (i.e., for any distribution $\mathcal{P}\in\mathcal{H}_{A}$). By the definition of $E_{j}$ being an e-value for each $j\in A$, for any $A\subseteq[K]$, we have that
$\mathbb{E}_{A}[E_{A}]\leq 1.$
We define the set of candidate discovery sets as
$\overline{\mathcal{C}}\coloneqq\left\{R\subseteq[K]:E_{A}\geq\frac{\textnormal{FDP}_{A}(R)}{\alpha}\quad\forall{A\subseteq[K]}\right\}.$ (3)
Proposition 1.
For any procedure $\mathcal{R}$ that ensures $R\in\overline{\mathcal{C}}$, we have that $\textnormal{FDR}(\mathcal{R})\leq\alpha$.
Proof.
For any $R\in\overline{\mathcal{C}}$, we have that
$\textnormal{FDP}_{A^{*}}(R)\leq\alpha E_{A^{*}}.$ (4)
Taking expectations, the claim is proved. ∎
Let $R_{i}\subseteq[K]$ be the set of hypotheses corresponding to the $i$ largest e-values (with ties being arbitrarily broken) for each $i\in[K]$. Define the procedure that outputs largest discovery set in $\overline{\mathcal{C}}$ as the ${\overline{\text{eBH}}}$ procedure:
$k^{\overline{\textnormal{eBH}}}$ $\coloneqq\max\ \{i\in[K]:R_{i}\subseteq\overline{\mathcal{C}}\}$ (5)
$R^{\overline{\textnormal{eBH}}}$ $\coloneqq\mathcal{R}^{\overline{\textnormal{eBH}}}(\mathbf{E})\coloneqq R_{k^{\overline{\textnormal{eBH}}}}.$ (6)
Note that $R^{\overline{\textnormal{eBH}}}\in\underset{R\subseteq\overline{\mathcal{C}}}{\textnormal{argmax}}|R|$ by definition. We will now show that this method is provably (usually strictly) more powerful as eBH, the standard way to control the false discovery rate (FDR) in this setting. Another way of interpreting is eBH is to first define self-consistent candidate discovery sets $\mathcal{C}^{\textnormal{SC}}$, and then note that $R^{\textnormal{eBH}}$ is the largest such set. Formally,
$\overline{\mathcal{C}}^{\textnormal{SC}}$$\coloneqq\left\{R\subseteq[K]:\min_{i\in R}\ E_{i}\geq K/(\alpha|R|)\right\},$ (7)
$R^{\textnormal{eBH}}$ $=\underset{R\subseteq\mathcal{C}^{\textnormal{SC}}}{\textnormal{argmax}}\ |R|.$ (8)
Proposition 2.
$\mathcal{C}^{\textnormal{SC}}\subseteq\overline{\mathcal{C}}$, and consequently $R^{\textnormal{eBH}}\subseteq R^{{\overline{\textnormal{eBH}}}}$.
Proof.
Consider the following derivation for arbitrary $A$ and $R$:
$E_{A}$ $=|A|^{-1}\sum\limits_{i\in A}E_{i}$ (9)
$\geq|A|^{-1}|A\cap R|\min_{i\in A\cap R}E_{i}\geq|A|^{-1}|A\cap R|\min_{i\in R}E_{i}\geq K^{-1}|A\cap R|\min_{i\in R}E_{i}.$ (10)
Recall that $E_{A}$ is defined as the arithmetic mean e-value in (2). The first inequality is lower bounding each term in the sum and also dropping some positive terms. The second inequality follows by strictly expanding the set which we take a minimum over. The third inequality is because $K\geq|A|$.
Now, assume that $\textnormal{FDP}_{A}(R)>0$ and $R\in\mathcal{C}^{\textnormal{SC}}$ (otherwise $\textnormal{FDP}_{A}(R)/E_{A}\leq\alpha$ is trivially satisfied). We have the following result for any $R\in\mathcal{C}^{\textnormal{SC}}$:
$E_{A}\geq K^{-1}|A\cap R|\min_{i\in R}E_{i}\geq\frac{|A\cap R|}{\alpha|R|}=\frac{\textnormal{FDP}_{A}(R)}{\alpha}.$ (11)
The first inequality is by (10), and the second inequality is by the definition of $\mathcal{C}^{\textnormal{SC}}$. Thus, we get that
$\frac{\textnormal{FDP}_{A}(R)}{E_{A}}\leq\alpha$ (12)
for any $R\in\mathcal{C}^{\textnormal{SC}}$ by (11). Thus, we get that $\mathcal{C}^{\textnormal{SC}}\subseteq\overline{\mathcal{C}}$ and also that $R^{\textnormal{eBH}}\in\mathcal{C}$ Since $R^{\overline{\textnormal{eBH}}}\in\underset{R\in\overline{\mathcal{C}}}{\text{argmax}}\ |R|$, we have that $|R^{\textnormal{eBH}}|=k^{\textnormal{eBH}}\leq k^{{\overline{\textnormal{eBH}}}}=|R^{\overline{\textnormal{eBH}}}|$ and consequently $R^{\textnormal{eBH}}=R_{k^{\textnormal{eBH}}}\subseteq R_{k^{\overline{\textnormal{eBH}}}}=R^{\overline{\textnormal{eBH}}}$. Thus, we have shown our desired results. ∎
The looseness in the inequalities in (10) hints at $R^{\overline{\textnormal{eBH}}}$ being potentially more powerful than eBH, and we will observe that is indeed the case in our simulations.
Input: E-values $(E_{1},\dots,E_{K})$ and target FDR level $\alpha\in[0,1]$
Output: Closed eBH discovery set $R^{\overline{\textnormal{eBH}}}$
for $k=K,K-1,\dots,k^{\textnormal{eBH}}+1$ do
       (Attempt to reject hypotheses with $k$ largest e-values)
       for $r=1$ to $k$ do
             for $m=r$ to $r+K-k$ do
                   Let $S_{(r,k)}=\sum_{i=1}^{r}E_{(K-k+i)}$(sum of $r$ smallest e-values among top $k$).
                   Let $S_{(m-r)}=\sum_{i=1}^{m-r}E_{(i)}$(sum of $m-r$ smallest e-values overall).
                   Let $E_{(k,r,m)}=\frac{S_{(r,k)}+S_{(m-r)}}{m}$, and $\textnormal{FDP}_{(k,r)}=\frac{r}{k}$.
                   if $\textnormal{FDP}_{(k,r)}/\alpha>E_{(k,r,m)}$ then
                         continue to next $k$.
                   end if
             end for
       end for
      return $R^{\overline{\textnormal{eBH}}}=R_{k}$ (discovery set of $k$ largest e-values)
end for
return $R^{\overline{\textnormal{eBH}}}=R^{\textnormal{eBH}}$
Algorithm 1 $O(K^{3})$ dynamic programming algorithm for computing discoveries made by the ${\overline{\text{eBH}}}$ procedure. Let $E_{(i)}$ denote the $i$th smallest e-value for each $i\in[K]$.
Efficiently computing $R^{\overline{\textnormal{eBH}}}$
We use a dynamic programming approach to efficiently compute $R^{\overline{\textnormal{eBH}}}$ — our approach is summarized in Algorithm 1. First, we note that computing $R^{\overline{\textnormal{eBH}}}$ reduces to computing the discovery set size $k^{\overline{\textnormal{eBH}}}\in[K]$ and setting the discovery set as the hypotheses with the $k^{\overline{\textnormal{eBH}}}$ largest e-values. Recall that the ${\overline{\text{eBH}}}$ procedure finds the largest rejection set $R\subseteq[K]$ such that the following condition is satisfied for all subsets $A\subseteq[K]$:
$\widehat{\textnormal{FDR}}(R)\coloneqq\max_{A\subseteq 2^{[K]}}\ \frac{\textnormal{FDP}_{A}(R)}{E_{A}}\leq\alpha,$
where we let $/0=0$ 0.
The algorithm iteratively considers each $R_{k}$, which contains the hypotheses corresponding to the $k$ largest e-values, and determines whether $R_{k}\in\overline{\mathcal{C}}$ is true. The algorithm starts from $K$ and stops as soon as it is able to discover $R_{k}$ or reaches the eBH discovery set, i.e., at $k=k^{\textnormal{eBH}}$.
To do this efficiently, we note that for each $r\in[k]$, we can fix $r$ to be the number of false discoveries. Then, to maximize $\textnormal{FDP}_{A}(R)/E_{A}$, $A$ includes $r$ true nulls in $R_{k}$ with the smallest e-values in $R_{k}$ and we only need to maximize over $m\in\{r,\dots,r+K-k\}$, the number of true nulls outside of the rejection set, where for each $m$ we pick the $m-r$ smallest e-values.
Let $A_{(k,r,m)}$ be the indices corresponding to the $r$ smallest e-values in $R_{k}$ and the $m-r$ smallest e-values outside of $R_{k}$. Then, we can say that
$\widehat{\textnormal{FDR}}(R_{k})=\max_{r\in[k],m\in\{r,r+K-k\}}\ \frac{r/k}{E_{A_{(k,r,m)}}}.$
Thus, we ended up with $O(K^{3})$ computation, since for each $k\in\{|R^{\textnormal{eBH}}|+1,\dots,K\}$, we can compute $\widehat{\textnormal{FDR}}(R_{k})$ as a maximization over a $O(K^{2})$ sized table which also only take $O(K^{2})$ time to compute.
3 A closure principle for FDR control
Now, we will argue that any FDR controlling procedure (i.e., including procedures that take inputs that are not e-values) can be improved by or reinterpreted as a procedure based on deriving e-values for every intersection hypotheses $\mathcal{H}_{A}$ for each $A\subseteq[K]$. Let $\mathcal{R}$ be a FDR controlling procedure (with discovery set $R=\mathcal{R}(\mathbf{X})$) if it satisfies
$\sup_{A\subseteq 2^{[K]}}\mathbb{E}_{A}\left[\textnormal{FDP}_{A}(R)\right]\leq\alpha.$ (13)
In this section, we now let $E_{A}$ denote any e-value for the intersection hypothesis $\mathcal{H}_{A}$. Now, consider any e-closed procedure $\overline{\mathcal{R}}$, which is any procedure that produces a discovery set that satisfies $\overline{R}\in\overline{\mathcal{C}}$ for the resulting candidate set $\overline{\mathcal{C}}$ in (3).
Theorem 1.
Any e-closed procedure $\overline{\mathcal{R}}$ (with corresponding discovery set $\overline{R}=\overline{\mathcal{R}}(\mathbf{X})$) which uses
$E_{A}=\frac{\textnormal{FDP}_{A}(R)}{\alpha}.$ (14)
to form a e-collection $(E_{A})_{A\subseteq[K]}$ ensures that in $R\in\overline{\mathcal{C}}$. Further, the only resulting e-closed procedure $\overline{\mathcal{R}}$ is equivalent to $\mathcal{R}$, i.e., $\overline{R}=R$.
Thus, any FDR controlling procedure can be represented as a e-closed procedure.
Proof.
Note that $E_{A}$ is a valid e-value for $\mathcal{H}_{A}$, since $\mathbb{E}_{A}[\textnormal{FDP}_{A}(R)]\leq\alpha$ by $R$ being a FDR controlling procedure. Furthermore, we can see that
$\max_{A\subseteq 2^{[K]}}\frac{\textnormal{FDP}_{A}(R)}{E_{A}}=\max_{A\subseteq[K]}\frac{\textnormal{FDP}_{A}(R)}{\textnormal{FDP}_{A}(R)/\alpha}=\alpha,$ (15)
when we treat $0/0=1$ in the above division. Thus, $R\in\overline{\mathcal{C}}$. As a result, there exists a e-closed procedure which can always choose $\overline{R}\in\overline{\mathcal{C}}$. On the other hand, for every $R^{\prime}\subseteq[K]$ with $|R^{\prime}|>|R|$, we have any $A=R\cap R^{\prime}\cup A^{\prime}$ where $\emptyset\subset A^{\prime}\subseteq R^{\prime}\setminus R$ has $E_{A}=\textnormal{FDP}_{A}(R)/\alpha<\textnormal{FDP}_{A}(R^{\prime})/\alpha$. Similarly, if $|R^{\prime}|<|R|$, we can choose $A^{\prime}\in R\setminus R^{\prime}$ instead and note the same inequality. Thus, $R$ is the only set in $\overline{\mathcal{C}}$.
Remark 1.
The offline multiple testing problem can be viewed as a composite generalized Neyman-Pearson (GNP) problem (Grünwald, 2024). In particular, $R$ is the maximally compatible decision rule for the multiple testing problem w.r.t. the e-collection derived (from $R$) in Theorem 1. However, admissibility results for maximally compatible decision rules from Grünwald (2024) do not directly apply to our setting since their assumption on action “richness” (continuity of loss — in our case, FDP — w.r.t. the action space) is violated in multiple testing, as there are only a finite set of values that the FDP can take.
4 The compound ${\overline{\text{eBH}}}$ procedure
Ignatiadis et al. (2025) showed that eBH was a universal procedure for FDR control, i.e., every FDR controlling procedure can be written as the eBH procedure applied to some set of compound e-values. For compound e-values, we let our data be $\mathbf{X}=\widetilde{\mathbf{E}}\coloneqq(\widetilde{E}_{1},\dots,\widetilde{E}_{K})$. Compound e-values are all nonnegative and satisfy the following condition
$\sum_{i\in A}\mathbb{E}_{A}[\widetilde{E}_{i}]\leq K.$ (16)
for any $A\subseteq[K]$. In words, no matter which $A$ constitutes the true set of null hypotheses $A^{*}$, the expected sum of the corresponding compound e-values cannot exceed $K$.
Our results are complementary to the above universality result on eBH. Instead of reducing every FDR procedure to using eBH with compound e-values, we instead reduce it the ${\overline{\text{eBH}}}$ procedure with an e-value for each intersection hypothesis. These two interpretations are interconnected, and we can formulate eBH applied to compound e-values as the application of the ${\overline{\text{eBH}}}$ procedure to an e-collection for the intersection hypotheses.
For any $A\subseteq[K]$, note that the (16) implies that
$E_{A}=K^{-1}\sum_{i\in A}\widetilde{E}_{i}$ (17)
is a valid e-value for $\mathcal{H}_{A}$. Let the compound ${\overline{\text{eBH}}}$ procedure be the ${\overline{\text{eBH}}}$ procedure that utilizes (17) for its e-collection instead of (2) and be denoted as $\tilde{R}^{{\overline{\textnormal{eBH}}}}\coloneqq\mathcal{R}^{{\overline{\textnormal{eBH}}}}(\widetilde{\mathbf{E}})$. Let $\tilde{R}^{\textnormal{eBH}}\coloneqq\mathcal{R}^{\textnormal{eBH}}(\widetilde{\mathbf{E}})$ now be the discovery set that results from applying the eBH procedure.
Proposition 3.
We have that $\tilde{R}^{\textnormal{eBH}}\subseteq\tilde{R}^{{\overline{\textnormal{eBH}}}}$.
Proof.
The proof is nearly identical to that of the proof of Proposition 2, albeit using the e-values defined in (17).
$E_{A}=K^{-1}\sum\limits_{i\in A}\widetilde{E}_{i}\geq K^{-1}|A\cap\tilde{R}|\min_{i\in A\cap\tilde{R}}\widetilde{E}_{i}\geq K^{-1}|A\cap\tilde{R}|\min_{i\in\tilde{R}}\widetilde{E}_{i}.$ (18)
The first inequality is lower bounding each term in the sum and also dropping some positive terms. The second inequality is by strictly expanding the set which we take a minimum over.
We have the following result for $\tilde{R}^{\textnormal{eBH}}$:
$E_{A}\geq K^{-1}|A\cap\tilde{R}^{\textnormal{eBH}}|\min_{i\in\tilde{R}^{\textnormal{eBH}}}\widetilde{E}_{i}.\geq\frac{|A\cap\tilde{R}^{\textnormal{eBH}}|}{\alpha|\tilde{R}^{\textnormal{eBH}}|}=\frac{\textnormal{FDP}_{A}(\tilde{R}^{\textnormal{eBH}})}{\alpha}.$ (19)
The first inequality is by (18), and the second inequality is by the definition of $\mathcal{R}^{\textnormal{eBH}}$. (19) implies that $\tilde{R}^{\textnormal{eBH}}\in\overline{\mathcal{C}}$, and consequently we get that $\tilde{R}^{\textnormal{eBH}}=R_{k^{\textnormal{eBH}}}\subseteq R_{k^{\overline{\textnormal{eBH}}}}=\tilde{R}^{\overline{\textnormal{eBH}}}$. Thus, we have shown our desired result.
We conjecture that the compound ${\overline{\text{eBH}}}$ procedure and standard compound eBH procedure are actually equivalent, as we have observed simulations where this is the case, and we will leave a rigorous investigation to future work.
5 Extensions
We discuss some extensions to the ${\overline{\text{eBH}}}$ procedure. These extensions primarily highlight the fact that the advantages of e-value procedures are retained for our procedure as well.
Post-hoc validity.
E-values have a notion of post-hoc validity that allows one to derive guarantees for a data-dependent choice of $\alpha$ (Grünwald, 2024; Koning, 2024), which the eBH procedure also possesses. This property is preserved for ${\overline{\text{eBH}}}$ as well.
Proposition 4.
The ${\overline{\text{eBH}}}$ procedure is valid for post-hoc (data-dependent) choice of $\alpha\in(0,1]$ in the following sense:
$\mathbb{E}\left[\frac{\textnormal{FDP}_{A^{*}}(R^{{\overline{\textnormal{eBH}}}_{\alpha}})}{\alpha}\right]\leq 1,$ (20)
where $R^{{\overline{\textnormal{eBH}}}_{\alpha}}$ is the ${\overline{\text{eBH}}}$ procedure applied at level $\alpha$.
Proof.
By construction, we have the following deterministic inequality
$\frac{\textnormal{FDP}_{A^{*}}(R^{{\overline{\textnormal{eBH}}}_{\alpha}})}{\alpha}\leq E_{A^{*}}$ (21)
that is satisfied even for possibly data-dependent $\alpha$. We get our desired result by taking expectations on both sides. ∎
Randomization.
Xu and Ramdas (2023) introduced stochastic rounding of e-values to use randomization to improve e-value based hypothesis tests as well as the eBH procedure. We can apply stochastic rounding individually to each intersection e-value $E_{A}$ for $A\subseteq[K]$, as follows.
Recall that $R_{k}$ denotes the set of hypotheses corresponding to the $k$ largest e-values, and solve for $\hat{k}:=k^{\overline{\textnormal{eBH}}}$ from (5). Then, define data-dependent test levels
$\hat{\alpha}_{A,k}\coloneqq\left(\frac{\textnormal{FDP}_{A}(R_{k})}{\alpha}\right)^{-1}$ (22)
for each $A\subseteq[K],k\in[K]$. Finally, define
$S_{A}(E_{A})$ =$=\begin{cases}\hat{\alpha}_{A,\hat{k}}^{-1}&\text{ if }\hat{\alpha}_{A,\hat{k}+1}>\hat{\alpha}_{A,\hat{k}}\text{ or }U>\frac{E_{A}-\hat{\alpha}^{-1}_{A,\hat{k}}}{\hat{\alpha}^{-1}_{A,\hat{k}+1}-\hat{\alpha}^{-1}_{A,\hat{k}}}\\ \hat{\alpha}_{A,\hat{k}+1}^{-1}&\text{ otherwise}\end{cases}.$ (23)
One can then check that $S_{A}(E_{A})$ is an e-value where $U$ is uniform random variable on $[0,1]$ that is independent of everything else.
Note that $S_{A}(E_{A})\geq\hat{\alpha}^{-1}_{A,k^{\overline{\textnormal{eBH}}}}$ almost surely, which means that ${\overline{\text{eBH}}}$ using the e-collection $(S_{A}(E_{A}))_{A\subseteq[K]}$ will never make fewer discoveries. Further, if $\min_{A\subseteq[K]}E_{A}-\hat{\alpha}^{-1}_{A,k^{{\overline{\textnormal{eBH}}}}}>0$, there is positive probability that $k^{\overline{\textnormal{eBH}}}+1$ discoveries can be made. In such a case, the randomized ${\overline{\text{eBH}}}$ procedure is strictly more powerful than the ${\overline{\text{eBH}}}$ procedure.
Boosting.
Wang and Ramdas (2022) observed that one can increase the power of an e-value when testing against a finite set of rejection thresholds. In the most general setting of assuming arbitrary dependence, if one knows the marginal distribution of an e-value under the null, one can calculate a boosting factor $b$ for an e-value $E$ as follows
$b\coloneqq\max\ \{b^{\prime}\in[1,\infty):\mathbb{E}[T(b^{\prime}E)]\leq 1\}$
where the expectation is taken under the known null distribution. Here, $T$ is a truncation function that takes the largest element in a predetermined set that is less than or equal to the input.
In our setting, one simple way to boost e-values is to define a different truncation function $T_{A}$ for each $A\subseteq[K]$ as follows:
$T_{A}(x)\coloneqq\max\ \{r/(\alpha k):r/(\alpha k)\leq x,k\in[K],r\in[k\wedge|A|]\}\cup\{0\},$
where $a\wedge b$ is the minimum of $a$ and $b$. This is a novel truncation function unlike ones used in boosting for the eBH procedure, and is derived from restricting the support of $E_{A}$ to only the values $\textnormal{FDP}_{A}(R)/\alpha$ can take. Let $b_{A}$ be the boosting factor for $E_{A}$ with truncation function $T_{A}$ for each $A\subseteq[K]$. Then, one can notice that $T_{A}(b_{A}E_{A})$ is a valid e-value and is at least as powerful as $E_{A}$ in the sense that $\textnormal{FDP}_{A}(R)/\alpha\leq E_{A}$ implies $\textnormal{FDP}_{A}(R)/\alpha\leq T_{A}(b_{A}E_{A})$ for all $R\subseteq[K]$.
Using the above boosted e-values immediately yields that the boosted ${\overline{\text{eBH}}}$ procedure that dominates the original. Other boosting techniques may be possible given more dependence assumptions or knowledge about the joint distribution of the entire e-collection $(E_{A})_{A\subseteq[K]}$.
Alternative e-merging functions.
While we used the average e-merging function in our formulation of the ${\overline{\text{eBH}}}$ procedure, one can use other e-merging functions when stronger dependence assumptions, i.e., independence or sequential dependence, are made about the e-values. For example, if we assume all null e-values are independent, one can let the e-values for each intersection hypothesis be the product of the base hypothesis e-values, i.e.,
$E_{A}=\prod\limits_{i\in A}E_{i}.$ (24)
One can then proceed to define a ${\overline{\text{eBH}}}$ procedure for the product e-merging function in a similar fashion as we did in Section 2, along with a corresponding algorithm, by using the logarithm to turn the products into sums when computing e-values for the intersection hypotheses.
6 Numerical simulations
We perform simulations for independent Gaussians where we let $\pi_{0}\coloneqq|A^{*}|/K$ be the null proportion, $\mu$ be the signal of the alternatives and let $X_{i}\sim\mathcal{N}(0,1)$ for $i\in A^{*}$ and $X_{i}\sim\mathcal{N}(\mu,1)$ otherwise for $i\in[K]\setminus A^{*}$. We let our e-value be $E_{i}\coloneqq\exp(\lambda|X_{i}|-\lambda^{2}/2)/2$ for $\lambda=\sqrt{\log(2K/\alpha)}$. We let $\alpha=0.1$ in our simulations. We plot the empirical average FDR and TDP $\coloneqq\mathbb{E}[|A^{*}\cap R|/|A^{*}|]$ of eBH, minimally adaptive eBH (Ignatiadis et al., 2024), and ${\overline{\text{eBH}}}$ over $n=1000$ trials. We plot the results in Figure 1, and can see that the ${\overline{\text{eBH}}}$ procedure improves noticeably over the standard and minimally adaptive eBH procedures across all settings, while controlling the FDR below $\alpha$.
Figure 1: Plot of empirical FDR and TDP for different null proportions $\pi_{0}$ and alternative mean $\mu=3$. All differences are significant since the closed eBH ($\overline{\textnormal{eBH}}$) procedure dominates the eBH and minimally adaptive eBH (eBH min-ada) procedures.
7 Related work
We will briefly describe the two lines of research and prior art that are most relevant to our current work. Aside from these, there is a substantial literature using e-values and the eBH procedure for FDR control — Ramdas and Wang (2025) performs a thorough overview of such developments.
Closed testing and e-values.
There is a vast amount of literature for FWER control and simultaneous (over all potential discovery sets) FDP bounds using closed testing — see overviews presented in Cui et al. (2021, Chapters 9 and 14) and Henning and and Westfall (2015). We will point out two specific works that are particularly relevant to ours since they also utilize e-values in the context of closed testing. The first is Vovk and Wang (2023), who formulate a method for FWER control via the arithmetic mean e-merging function, and consequently also simultaneous FDP bounds. Hartog and Lei (2025) extend the benefit of using arithmetic mean e-merging functions to graphical approaches for FWER control (Bretz et al., 2009), and derive computationally efficient algorithms for e-value versions of the Fallback (Wiens and and Dmitrienko, 2005) and Holm (1979) procedures. Our work considers the FDR metric (rather than FWER or probabilistic bounds on the FDP) which has not been connected to intersection hypothesis testing or e-merging functions before.
Boosting the power of eBH
Another line of work that is pertinent is various approaches to boosting the eBH procedure so that more discoveries are made. There are generally two areas in this line: (1) boosting with no assumptions on the e-value distributions and (2) boosting e-values by using explicit knowledge about the underlying (marginal or joint) distribution.
For the first area, we have already discussed the connection between our work and the stochastic rounding framework of Xu and Ramdas (2023) in Section 5. Another particularly relevant work is that of Ignatiadis et al. (2024, Section S3.2), which introduces a minimally adaptive eBH procedure where one first tests the global null with the the arithmetic mean of the e-values, and only once the global null rejects, the eBH procedure is applied at level $\alpha K/(K-1)$. They show that this method dominates eBH as well. This method is, in turn, dominated by the ${\overline{\text{eBH}}}$ procedure since rejection of the global null implies that $E_{[K]}\geq\alpha^{-1}\geq\textnormal{FDP}(R)/\alpha$ is satisfied for all $R\subseteq[K]$ (where $E_{[K]}$ is the arithmetic mean of all e-values as defined in (2)). One can then show that the $R$ which results from eBH applied at a level of $\alpha K/(K-1)$ implies that $R\in\overline{\mathcal{C}}$ using an argument similar to that in the proof of Proposition 2. We see the empirical implications of this in our simulations in Section 6.
The second area includes two general types of assumptions.
  • Marginal assumptions: Along with introducing the eBH procedure, Wang and Ramdas (2022) also outlined how one can boost e-values if one knows the explicit marginal distribution of each e-value, and we discuss a simple variant for boosting the ${\overline{\text{eBH}}}$ procedure in Section 5. Blier-Wong and Wang (2024) relaxed the necessary knowledge on marginal distributions for boosting by deriving boosting thresholds for certain nonparametric classes of e-values.
  • Conditional assumptions: Lee and Ren (2024); Lee et al. (2025) improve the power of eBH in the setting where one knows the distribution of all $K$ e-values when conditioned a sufficient statistic calculated from the data under a singular null (i.e., a specific hypothesis is null, and all other hypotheses may be either null or non-null). This can also be viewed as a framework for deriving compound e-values from e-values when their joint distribution is known. Boosting factors can be calculated for each e-value via exact knowledge of these conditional distributions.
Developments in this area are orthogonal to our contribution. These boosting methods all require additional knowledge about the underlying distributions. On the other hand, the ${\overline{\text{eBH}}}$ procedure requires no assumptions on the e-values, and can be applied with valid FDR control in all situations where the eBH procedure can be applied. Potentially, these boosting methods can be used in conjunction with ${\overline{\text{eBH}}}$ to further improve its power.
8 Conclusion
We have shown in this manuscript that the eBH procedure is inadmissible, and can be improved through a FDR closure principle that results in the ${\overline{\text{eBH}}}$ procedure. We provide a computationally efficient way of computing the ${\overline{\text{eBH}}}$ discovery set. Further, our FDR closure principle dictates that all FDR controlling procedures rely on e-values for testing each possible intersection hypothesis of the base hypotheses. These results open many lines of future work, particularly in practical and computationally efficient applications of the closure principle to improving existing FDR controlling procedures.